# Kernel Three Pass Regression Filter

### v3

@@ 6.2 Results
0 12.36 0.28 1.28 2.27 5.54 11.79 66.48
5 13.07 0.85 1.99 2.41 6.11 13.21 68.32
10 14.06 1.28 2.56 2.84 7.24 16.90 70.60
20 17.05 2.98 3.98 5.11 9.94 26.56 76.42
Excluding AR
0 - 1.42 1.56 2.84 5.47 13.00 75.71
5 - 2.84 2.06 4.76 5.75 17.97 78.76
10 - 5.26 3.27 7.74 7.03 25.99 81.53
20 - 11.08 5.89 14.35 11.43 41.34 86.08
Table 10: Distribution of Best Forecasting Methods Across All Series (Percentage)
We present the results in table-10. The findings presented above yield several noteworthy observations. First, it is evident that unsupervised forecasting techniques, including PCR, Squared-PC, PC-squared, and kernel PCA, exhibit inferior performance across the majority of scenarios when compared to our method. Second, our method, kernel 3PRF, demonstrates unequivocal superiority in longer-horizon forecasting endeavors. Third, our method is unequivocally superior across all horizons when autoregressive (AR) method is excluded. Our method does not outperform AR in the short term, but its performance remains competitive, often closely trailing the best short-run autoregressive method. This can be seen by increasing the tolerance level. The instances where our method can be labeled as ‘best’ increase rapidly as we increase the tolerance level.
@@ 7 Conclusion
Building upon the three-pass regression filter by Kelly & Pruitt, (2015), we introduce a new forecasting method, kernel three-pass regression filter. Through extensive empirical exercises, we show that this approach holds promise as a dependable forecasting tool. Improved performance can be attributed to two noteworthy features of our method. First, it integrates non-linear relationships by transforming input data into a higher-dimensional space, encapsulating its non-linear functions. Second, it operates as a supervised method, effectively filtering out and discarding irrelevant factors while predicting the target variable.
@@ A.1 Proofs of Theoretical Results
Proof: Proof can be seen from Kelly & Pruitt, (2015), Lemma 2 in their appendix. The only difference is the omission of the matrix $\boldsymbol{J}_{N}$ in the various expressions. This, however, doesn’t affect the rates, as can be verified from their proofs. We do not allow an intercept in pass-2 because doing so will require demeaning of the transformed predictor(s), which is not feasible.
@@
    $M^{-1}T^{-2}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{y}$ $=\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\beta}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$
@@
    $M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ $=\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$
@@
$+\boldsymbol{\Lambda}\left(M^{-1}T^{-1}\boldsymbol{F}^{\prime}\boldsymbol{J}_{T}\boldsymbol{\varepsilon}\boldsymbol{\Phi}\right)\boldsymbol{F}_{t}+\boldsymbol{\Lambda}\left(M^{-1}T^{-1}\boldsymbol{F}^{\prime}\boldsymbol{J}_{T}\boldsymbol{\varepsilon}\varepsilon_{t}\right)$
@@
$+\left(M^{-1}T^{-1}\boldsymbol{\omega}^{\prime}\boldsymbol{J}_{T}\boldsymbol{\varepsilon}\boldsymbol{\Phi}\right)\boldsymbol{F}_{t}+\left(M^{-1}T^{-1}\boldsymbol{\omega}^{\prime}\boldsymbol{J}_{T}\varepsilon\boldsymbol{\varepsilon}_{t}\right)$
@@
Item 3: Let = $\hat{\boldsymbol{F}}_{C,t}=M^{-1}T^{-1}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{x}_{t})$. Then, given Lemma 2.1, standard arguments would imply that $M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ $=\dfrac{\hat{\boldsymbol{F}}_{C}\boldsymbol{J}_{T}{\hat{\boldsymbol{F}}_{C}}^{\prime}}{T}$
$=\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\left(T^{-1}\boldsymbol{F}\boldsymbol{J}_{T}\boldsymbol{F}\right)\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$. Given Assumption 2.1, we have that
$\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\left(T^{-1}\boldsymbol{F}\boldsymbol{J}_{T}\boldsymbol{F}\right)\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}$ $=\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left({T}^{-1/2}\right)$.
Therefore, we have that, $M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ $=\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)+\boldsymbol{O}_{p}\left({T}^{-1/2}\right)=$ $\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$.
@@
where $\boldsymbol{H}_{f}\equiv\hat{\boldsymbol{F}}_{A}\hat{\boldsymbol{F}}_{B}^{-1}\mathbf{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}$
@@
The second equality follows from Lemma 2.1 and the final equality uses the definition of $\boldsymbol{H}_{f}$.
@@
$\hat{\boldsymbol{\beta}}-\boldsymbol{G}_{\beta}\boldsymbol{\beta}=\boldsymbol{O}_{p}(\delta_{MT}^{-1}).$
where $\boldsymbol{G}_{\beta}\equiv\hat{\boldsymbol{\beta}}_{1}^{-1}{\hat{\boldsymbol{\beta}}_{2}}\left[\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\right]^{-1}\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}$,
$\hat{\boldsymbol{\beta}}_{1}=\hat{\boldsymbol{F}}_{A}$ and $\hat{\boldsymbol{\beta}}_{2}=\hat{\boldsymbol{F}}_{B}$
@@
×$\times\left(M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\right)^{-1}M^{-1}T^{-2}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{y}$
@@
$\left(\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\right)^{-1}-\left(\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\right)^{-1}\boldsymbol{O}_{p}(\delta_{MT}^{-1})\left(\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})\right)^{-1}.$ The last equality uses the definition of $\boldsymbol{G}_{\beta}$.
@@
$=\beta_{0}+\bar{\boldsymbol{f}}^{\prime}\boldsymbol{\beta}_{f}+O_{p}(T^{-1/2})+\left(\boldsymbol{H}_{f}\boldsymbol{f}_{t}+\boldsymbol{O}_{p}(\Gamma_{NT}^{-1})\right)^{\prime}\left(\boldsymbol{G}_{\beta}\boldsymbol{\beta}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})\right)$
$=\beta_{0}+\bar{\boldsymbol{f}}^{\prime}\boldsymbol{\beta}_{f}+\left(\boldsymbol{f}_{t}-\bar{\boldsymbol{f}}\right)^{\prime}\boldsymbol{H}_{f}^{\prime}\boldsymbol{G}_{\beta}\boldsymbol{\beta}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})$
$=\beta_{0}+\bar{\boldsymbol{f}}^{\prime}\boldsymbol{\beta}_{f}+\left(\boldsymbol{f}_{t}-\bar{\boldsymbol{f}}\right)^{\prime}\boldsymbol{\beta}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})$
$=\beta_{0}+\boldsymbol{f}_{t}^{\prime}\boldsymbol{\beta}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})$
@@
The second equality follows from lemma 1.15. The fourth equality follows if $\boldsymbol{H}_{f}^{\prime}\boldsymbol{G}_{\beta}$ is an identity matrix. This is indeed true since $\boldsymbol{H}_{f}^{\prime}\boldsymbol{G}_{\beta}$ = $\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\left[\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\right]^{-1}\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}$. Using the arguments as in Lemma 5 and Theorem 1 of Kelly & Pruitt, (2015) the RHS is an identity matrix, given assumptions 4 and 5.

### v4

@@ 6.2 Results
0 4.83 7.10 0.14 1.28 1.85 5.68 0.43 11.51 67.19
5 7.10 10.23 0.71 1.70 1.85 6.11 0.57 13.21 69.18
10 9.80 11.65 1.28 2.27 2.27 7.39 0.71 17.19 72.16
20 13.64 16.34 2.70 3.69 4.69 10.37 1.14 27.70 77.84
Without AR
0 - - 1.21 1.42 2.84 5.47 0.71 12.86 75.50
5 - - 2.70 1.99 4.62 5.68 0.85 17.83 78.55
10 - - 5.11 3.20 7.53 6.96 1.07 25.99 81.39
20 - - 11.01 5.75 14.20 11.43 1.49 41.34 85.94
Notes: We call a method the best if either it offers maximum out of sample (OOS) $R^{2}$ or its OOS $R^{2}$ falls within a tolerance percentage of the highest OOS $R^{2}$ for a given row. We run this experiment on all 176 quarterly time series available in the FRED-QD dataset for 8 forecast horizons i.e. a total of $176\times 8=1408$ comparisons.
We present the results in table-10. The findings presented above yield several noteworthy observations. First, it is evident that unsupervised forecasting techniques, including PC, Squared-PC, PC-squared, Autoencoder, and kernel PCA, exhibit inferior performance across the majority of scenarios when compared to our method. Second, our method, kernel 3PRF, demonstrates unequivocal superiority in longer-horizon forecasting endeavors. Third, our method is unequivocally superior across all horizons when autoregressive (AR) and closely related DI methods are excluded. Fourth, even in the short run, our method is the single best-performing approach under strict (zero-tolerance) evaluations of absolute performance. Moreover, its performance remains comparable to the leading short-run benchmarks, AR(p) and DI, when we allow for some tolerance level.
Taken together, by demonstrating consistent performance across all series in our dataset, we conclude that our method can be regarded as a reliable forecasting tool across all forecast horizons.
@@ 7 Conclusion
Building upon the three-pass regression filter by kelly, we introduce a new forecasting method, kernel three-pass regression filter. Through extensive empirical exercises, we show that this approach holds promise as a dependable forecasting tool. Improved performance can be attributed to two noteworthy features of our method. First, it integrates non-linear relationships by transforming input data into a higher-dimensional space, encapsulating its non-linear functions. Second, it operates as a supervised method, effectively filtering out and discarding irrelevant factors while predicting the target variable.
@@ A.1 Proofs of Theoretical Results
Proof: The result follows from kelly, Lemma 2 in their appendix. In their setup, the matrix $\boldsymbol{J}_{N}$ appears in these expressions, whereas in our case there is no analogous $\boldsymbol{J}_{M}$. Consequently, if we expand the matrix products as in kelly, fewer terms arise compared to their setting. The order of the remaining terms was established in kelly, with the only adjustment here being that we work with $\varphi(\boldsymbol{X})$ instead of $\boldsymbol{X}$, so the feature dimension $M$ appears in the relevant norms and sums. Therefore, the convergence rates of these matrix products are at least as fast, if not faster, than analogous products in kelly.
Assumptions 3.5-3.7 in our paper effectively replace Assumption 4 of kelly, which imposes a distributional structure on these stochastic processes. For the results featuring in Lemma 1 (subsequently used in Lemma 2) of kelly, they only use the boundedness of second moments of these processes, which is precisely what we assume in 3.5-3.7. Therefore, the convergence results here remain valid. The stronger distributional assumptions in kelly are used only to establish convergence in distribution of the factors, coefficients, and target, which we do not pursue here due to the concerns flagged in the remarks of our paper.
@@
    $M^{-1}T^{-2}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{y}$ $\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\beta}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$
@@
    $M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ $\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$
@@
$+\boldsymbol{\Lambda}\left(M^{-1}T^{-1}\boldsymbol{F}^{\prime}\boldsymbol{J}_{T}\boldsymbol{\varepsilon}\boldsymbol{\Phi}\right)\boldsymbol{F}_{t}+\boldsymbol{\Lambda}\left(M^{-1}T^{-1}\boldsymbol{F}^{\prime}\boldsymbol{J}_{T}\boldsymbol{\varepsilon}\boldsymbol{\varepsilon}_{t}\right)$
@@
$+\left(M^{-1}T^{-1}\boldsymbol{\omega}^{\prime}\boldsymbol{J}_{T}\boldsymbol{\varepsilon}\boldsymbol{\Phi}\right)\boldsymbol{F}_{t}+\left(M^{-1}T^{-1}\boldsymbol{\omega}^{\prime}\boldsymbol{J}_{T}\boldsymbol{\varepsilon}\boldsymbol{\varepsilon}_{t}\right)$
@@
Item 3: Let = $\hat{\boldsymbol{F}}_{C,t}=M^{-1}T^{-1}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{x}_{t})$. Then, given Lemma 2.1, standard arguments would imply that $M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ $\dfrac{\hat{\boldsymbol{F}}_{C}\boldsymbol{J}_{T}{\hat{\boldsymbol{F}}_{C}}^{\prime}}{T}$
$\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\left(T^{-1}\boldsymbol{F}\boldsymbol{J}_{T}\boldsymbol{F}\right)\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$. Given Assumption 2.1, we have that
$\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\left(T^{-1}\boldsymbol{F}\boldsymbol{J}_{T}\boldsymbol{F}\right)\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}$ $\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left({T}^{-1/2}\right)$.
Therefore, we have that, $M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}$ $\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)+\boldsymbol{O}_{p}\left({T}^{-1/2}\right)=$ $\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$.
@@
where $\boldsymbol{H}_{f}\equiv\hat{\boldsymbol{F}}_{A}\hat{\boldsymbol{F}}_{B}^{-1}\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}$
@@
$=\hat{\boldsymbol{F}}_{A}\hat{\boldsymbol{F}}_{B}^{-1}\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}\boldsymbol{f}_{t}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$
@@
The second equality follows from Lemma 2.1. The second last equality follows from the same reasoning as in the Proof of Theorem 1 in kelly, noting that $\mathcal{P}_{1}$ as described in their setup is an identity matrix by Assumption 4. Final equality uses the definition of $\boldsymbol{H}_{f}$.
@@
$\hat{\boldsymbol{\beta}}-\boldsymbol{G}_{\beta}\boldsymbol{\beta}_{f}=\boldsymbol{O}_{p}(\delta_{MT}^{-1}).$
where $\boldsymbol{G}_{\beta}\equiv\hat{\boldsymbol{\beta}}_{1}^{-1}{\hat{\boldsymbol{\beta}}_{2}}\left[\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}^{3}\boldsymbol{\Lambda}_{f}^{\prime}\right]^{-1}\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}^{2}$, where
$\hat{\boldsymbol{\beta}}_{1}=\hat{\boldsymbol{F}}_{A}$ and $\hat{\boldsymbol{\beta}}_{2}=\hat{\boldsymbol{F}}_{B}$, and $\boldsymbol{\Delta}_{f}$ is defined in the paragraph preceding Theorem 1. ,
@@
$\times\left(M^{-2}T^{-3}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{Z}\right)^{-1}M^{-1}T^{-2}\boldsymbol{Z}^{\prime}\boldsymbol{J}_{T}\varphi(\boldsymbol{X})\varphi(\boldsymbol{X})^{\prime}\boldsymbol{J}_{T}\boldsymbol{y}$
@@
$=\hat{\boldsymbol{\beta}}_{1}^{-1}{\hat{\boldsymbol{\beta}}_{2}}\left[\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}^{3}\boldsymbol{\Lambda}_{f}^{\prime}\right]^{-1}\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{F}^{2}{\boldsymbol{\beta}}_{f}+\boldsymbol{O}_{p}\left(\delta_{MT}^{-1}\right)$
@@
$\left(\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\right)^{-1}-\left(\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}\right)^{-1}\boldsymbol{O}_{p}(\delta_{MT}^{-1})\left(\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})\right)^{-1}.$ The last equality uses the definition of $\boldsymbol{G}_{\beta}$. The second-to-last equality follows the logic in the proof of Theorem 1 of kelly. Under Assumption 5 the normalization assumption 4, we can partition
$\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}=\left[\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}\mathcal{P}_{f}\boldsymbol{\Delta}_{f},\ \mathbf{0}\right],$
where $\mathcal{P}_{f}$ is the covariance matrix of the relevant factors and equals the identity matrix under Assumption 4. Consequently,
$\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\Lambda}^{\prime}=\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}\mathcal{P}_{f}\boldsymbol{\Delta}_{f}\mathcal{P}_{f}\boldsymbol{\Delta}_{f}\boldsymbol{\Lambda}_{f}^{\prime}=\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}^{3}\boldsymbol{\Lambda}_{f}^{\prime},$
and similarly,
$\boldsymbol{\Lambda}\boldsymbol{\Delta}_{F}\mathcal{P}\boldsymbol{\Delta}_{F}\boldsymbol{\beta}=\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}\mathcal{P}_{f}\boldsymbol{\Delta}_{f}\boldsymbol{\beta}_{f}=\boldsymbol{\Lambda}_{f}\boldsymbol{\Delta}_{f}^{2}\boldsymbol{\beta}_{f}.$
@@
$=\beta_{0}+\bar{\boldsymbol{f}}^{\prime}\boldsymbol{\beta}_{f}+O_{p}(T^{-1/2})+\left(\boldsymbol{H}_{f}\boldsymbol{f}_{t}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})\right)^{\prime}\left(\boldsymbol{G}_{\beta}\boldsymbol{\beta}_{f}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})\right)$
$=\beta_{0}+\bar{\boldsymbol{f}}^{\prime}\boldsymbol{\beta}_{f}+\left(\boldsymbol{f}_{t}-\bar{\boldsymbol{f}}\right)^{\prime}\boldsymbol{H}_{f}^{\prime}\boldsymbol{G}_{\beta}\boldsymbol{\beta}_{f}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})$
$=\beta_{0}+\bar{\boldsymbol{f}}^{\prime}\boldsymbol{\beta}_{f}+\left(\boldsymbol{f}_{t}-\bar{\boldsymbol{f}}\right)^{\prime}\boldsymbol{\beta}_{f}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})$
$=\beta_{0}+\boldsymbol{f}_{t}^{\prime}\boldsymbol{\beta}_{f}+\boldsymbol{O}_{p}(\delta_{MT}^{-1})$
@@
The second equality follows from lemma 1.15. The fourth equality follows if $\boldsymbol{H}_{f}^{\prime}\boldsymbol{G}_{\beta}$ is an identity matrix.